40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 23 23:55:03 2010, Katharine, Sharmila, Rana, Steve and Kiwamu, Update, VAC, wiped the BS 
    Reply  Thu Jun 24 11:16:32 2010, Sharmila, Rana and Kiwamu, HowTo, VAC, Inspection of the BS (sorry, no sounds) 
Message ID: 3111     Entry time: Wed Jun 23 23:55:03 2010     Reply to this: 3114
Author: Katharine, Sharmila, Rana, Steve and Kiwamu 
Type: Update 
Category: VAC 
Subject: wiped the BS 

Some unused optics in the BS chamber were removed. 

After that the beam splitter has been drag wiped. 

So now the BS chamber is waiting for the installation of  the other core optics i.e. PRM, SRM and Tip-Tilts. 

 

-- removing of unused optics

      There were some unused optics, mainly 1.5 inch optics which had been used for the oplevs in the chamber. 

Kathaine, Shamila (Team Magnet) and Kiwamu took those optics out from the chamber.

And then we carefully wrapped each of them by aluminum foils and put them in some clear boxes.

In fact, before wrapping them, we gently attached lens papers on their HR surfaces such that aluminum foils can not damage it.

Now there are only three 1.5 inch optics in the chamber, and they are supposed to be used for the oplevs.

We didn't remove any of the 2 inch optics and the PZT mirrors because they are still going to be used.

These are the pictures of the BS chamber after we cleaned up them. 

 

-- wiping of the BS

        Rana and Kiwamu drag wiped the HR surface of the BS by using lens paper with the solvents.

The below is the procedure we did. You can find some details about the wiping technique for suspended optics in this entry.

In this time we could wipe the beam splitter without removing the front earthquake stops because the beam splitter was brought close enough to us. 

 

(1). put some blocks attaching the edge of the bottom plate of the tower in order to record the original position.

(2).  locked the beam splitter to the frame by screwing the earthquake stops.

(3). made sure if it is really locked by seeing the output signal of the OSEMs in dataviewer. If it's locked successfully, the resonant frequency gets higher and the Q-value gets lower.

(3).  moved the BS tower close to the door in order to reach the beam splitter easily.

(4). inspected the surface by using a fiber light. There were about 10 bright spots on the HR surface.

(5). wiped the surface three times by using the lens paper with Aceton.

(6). wiped it several times with Isopropyl.

(7). inspected the surface again, found there were no big bright spots near the center. Thumbed up 

(8). put the tower back to the original place and released the beam splitter from the earthquake stops.

ELOG V3.1.3-